What is the change of temperature (in oC) when 175 J of energy is applied to 35.4 g of silver (specific heat of silver = 0.240 J/goC)? Round your answer to three sig figs.
1
Expert's answer
2020-05-08T14:11:47-0400

Q=mcΔT.

ΔT=Q/mc=175/35.4 x 0.24=‭20,59792843691149‬ oC